# Structural-Spectral Graph Convolution with Evidential Edge Learning for Hyperspectral Image Clustering

### v1

@@ IV-G Further Analysis of Evidence Guided Adaptive Edge Learning
TABLE VII: Ablation Study Results of Evidence Guided Adaptive Edge Learning.
$w^{pre}$ $w^{emp}$     Indian Pines     Pavia University     Botswana     Trento
    ACC NMI ARI     ACC NMI ARI     ACC NMI ARI     ACC NMI ARI
@@
TABLE VIII: Classification Accuracy (%) of Correct and Incorrect Edges with Initial Edge Weights ($w^{init}$) and Predicted Edge Weights ($w^{pre}$).
@@
$w^{init}$     62.33 69.89 88.69 97.40
$w^{pre}$     68.97 73.54 90.59 97.64
Inc.$\uparrow$     6.64 3.65 1.90 0.24
To validate the effectiveness of $w^{pre}$ and $w^{emp}$, we first conducted an ablation study. As shown in Table VII, the complete EGAEL module results in an average ACC improvement of 3.12%, while leveraging either $w^{pre}$ or $w^{emp}$ individually yields inferior performance, underscoring the importance of their interaction within EGAEL module.
Then, we interpreted $w^{pre}$ as the probability of an edge being correct, and determined an optimal threshold to classify correct and incorrect edges. The classification performance is compared against that of using initial edge weights $w^{init}$, which assume all edges are correct. As shown in Table VIII, $w^{pre}$ achieves higher accuracy on all the four datasets, with an improvement of 6.64% on Indian Pines dataset in particular. This demonstrates that the adjacency matrix constructed from $w^{pre}$ is more reliable than the initial one, thereby validating the effectiveness of the proposed EGAEL module.

### v2

@@ IV-E Running Time Analysis
TABLE VI: Ablation Study Results of Evidence Guided Adaptive Edge Learning.
Items     Indian Pines     Pavia University     Botswana     Trento
$w^{pre}$ $w^{emp}$     ACC NMI ARI     ACC NMI ARI     ACC NMI ARI     ACC NMI ARI
@@ IV-F Further Analysis of Structural-Spectral Graph Convolutional Operator
According to the experimental results presented in Table V, the proposed SSGCO outperforms all other variant networks with a large margin. In addition, we have identified two observations.
IV-G Further Analysis of Evidence Guided Adaptive Edge Learning
To validate the effectiveness of $w^{pre}$ and $w^{emp}$, we first conducted an ablation study. As shown in Table VI, the complete EGAEL module results in an average ACC improvement of 3.12%, while leveraging either $w^{pre}$ or $w^{emp}$ individually yields inferior performance, underscoring the importance of their interaction within EGAEL module.
Then, we interpreted $w^{pre}$ as the probability of an edge being correct, and determined an optimal threshold to classify correct and incorrect edges. The classification performance is compared against that of using initial edge weights $w^{init}$, which assume all edges are correct. As shown in Table VII, $w^{pre}$ achieves higher accuracy on all the four datasets, with an improvement of 6.64% on Indian Pines dataset in particular. This demonstrates that the adjacency matrix constructed from $w^{pre}$ is more reliable than the initial one, thereby validating the effectiveness of the proposed EGAEL module.
TABLE VII: Classification Accuracy of Correct and Incorrect Edges with Initial Edge Weights ($w^{init}$) and Predicted Edge Weights ($w^{pre}$).
@@ IV-G Further Analysis of Evidence Guided Adaptive Edge Learning
$w^{init}$     62.33% 69.89% 88.69% 97.40%
$w^{pre}$     68.97% 73.54% 90.59% 97.64%
[…]
@@ V Conclusion
Nevertheless, the proposed method still has some limitations. At the beginning of our method, we represent each superpixel as the average of its internal pixel features. This common practice, however, leads to the loss of 2D spatial structure within superpixels. In future work, we aim to exploit superpixels’ internal spatial information. For example, we will consider padding superpixels into regular rectangles and introducing 2D convolution to learn better representations.